# A Large Deviations Perspective on Policy Gradient Algorithms

### v2

@@ Abstract
Motivated by policy gradient methods in the context of reinforcement learning, we derive the first large deviation rate function for the iterates generated by stochastic gradient descent for possibly non-convex objectives satisfying a Polyak-Łojasiewicz condition. Leveraging the contraction principle from large deviations theory, we illustrate the potential of this result by showing how convergence properties of policy gradient with a softmax parametrization and an entropy regularized objective can be naturally extended to a wide spectrum of other policy parametrizations.
Keywords— large deviations, reinforcement learning
2020 Mathematics Subject Classification— 60F10 $\cdot$ 90C26
@@
Policy gradient methods are at the core of several reinforcement learning (RL) algorithms. e.g., see [Sut+99, DPS12, Aga+21]. As such, a wide adoption of these algorithms is aided by precisely understanding their performance. To that end, in scenarios where policy gradients can only be evaluated stochastically, it is crucial to understand the underlying distribution governing the evolution of policy iterates. Despite the popularity of stochastic policy gradient methods for over 20 years, global convergence has been only understood recently [Faz+18, Zha+20, BR19, Aga+21]. Although global convergence proofs are actively developed for different policy gradient algorithm templates [BR19, Aga+21, Cen+22], two aspects of the current convergence analysis would benefit from further study: 1) global convergence guarantees are often stated in terms of expected suboptimality; and 2) the choice of policy parametrizations impacts convergence behavior [Mei+20]. In light of these observations, this paper aims to derive sharp convergence rates in probability and to provide a unifying approach towards understanding the effect of different policy parametrizations. As we build upon a rich history of work, we succinctly comment on: (i) high probability analysis in stochastic optimization and RL; and (ii) policy gradient methods in RL.
@@
(i) Despite stochastic gradient descent (SGD) being over 70 years old [RM51], there has been a surge of interest in terms of high probability analysis for (SGD): for strongly-convex objectives [Har+19]; non-convex objectives satisfying a Polyak-Łojasiewicz (PL) condition [MDB20]; and non-convex objectives [GL13, Liu+23] with Lipschitz gradients. A motivation for studying tight convergence bounds in terms of probabilities over convergence in expectation is due to the fact that the iterates generated by SGD can be brittle and large deviations from the expected value can occur, e.g., see [Gow+20] and references therein. Hence, a high-probability bound is attractive since it reassures the practitioner that certain behavior occurs at least with a close-to-$1$ probability for a single realization of iterates.
@@
(ii) Policy gradient methods can be interpreted as gradient descent applied to the policy optimization problem in RL [SB18, Ch. 13], see in particular [SB18, p. 337] for historical remarks. Therefore, when the state space is large [Bot10] and due to necessary approximation, one oftentimes resorts to stochastic variants, whose convergence guarantees are often stated in terms of expectations [Lan23], with the exception of [DZL21], see also [MDB20, § 2.2]. The high-probability convergence rate provided in [DZL21] can be seen as a loosened bound of the large deviation rate we aim to derive. The pursuit of a tight high-probability concentration bound for the iterations produced by stochastic policy gradient methods holds significant potential for bolstering the practical implementation and interpretability of policy gradient methods, thereby improving the overall applicability of RL. Studying the concentration behavior of SGD iterates through the lens of large deviations is pioneered by [BJK23]. It is worth mentioning that, however, their analysis exploits strong convexity of the optimization objective. Then, a first step towards understanding more precisely the concentration behavior of stochastic policy gradient iterates would be to study its entropy-regularized variant, which already faces the difficulties of non-convexity and non-uniformity of the PL constant.
@@
As the theory of large deviations is the key tool in this paper, we briefly introduce it already at this stage. Let $X_{t}\in\mathbb{R}^{d}$ be the $t^{\mathrm{th}}$ iterate of some stochastic algorithm aimed at driving $X_{t}$ to $x^{\star}$ for $t\to+\infty$, then, many aforementioned works provide statistical guarantees of the form $\mathbb{P}(\|X_{t}-x^{\star}\|_{2}\leq\varepsilon)\geq 1-\beta$ $\forall t\geq T$ for an appropriately chosen triple $(T,\varepsilon,\beta)$. Although often impressive pieces of work, there are some remarks to be made. First, it is often not clear if these high-probability bounds are tight. Secondly, the discrepancy is often measured by a sufficiently simple function, like the $\ell_{2}$-norm in this case, however, in practice one might be interested in vastly different sublevel sets, e.g., sublevel sets capturing harmful events, the ones that are of interest in safe RL. Third, the guarantees are oftentimes states for particular policy parametrizations, a unifying framework the generalizes performance analysis across different policy parametrizations remains largely unexplored. The theory of large deviations is precisely powerful when one aims to address these aforementioned points. Let us clarify terminology and provide intuition, yet, while skipping some details. A sequence of finite (probability) measures $(\mu_{t})_{t}$ on $\mathbb{R}^{d}$ is said to satisfy a large deviation principle (LDP) with rate function $I:\mathbb{R}^{d}\to[0,+\infty]$ when for any Borel set $\Theta\subseteq\mathbb{R}^{d}$
$-\overset{\circ}{r}:=-\inf_{\theta\in\overset{\circ}{\Theta}}I(\theta)\leq\liminf_{t\to+\infty}\frac{1}{t}\log\mu_{t}({\Theta})\leq\limsup_{t\to+\infty}\frac{1}{t}\log\mu_{t}({\Theta})\leq-\inf_{\theta\in\overline{\Theta}}I(\theta)=:-\overline{r}.$ (1.1)
@@
Now, identifying $(\mu_{t})_{t}$ with a sequence of random variables $(X_{t})_{t}$, we have after rearranging (1.1) that $\mathrm{exp}(-\overset{\circ}{r}t+o(t))\leq\mathbb{P}(X_{t}\in\Theta)\leq\mathrm{exp}(-\overline{r}t+o(t)).$ As such, an LDP captures—possibly tight, e.g., under mild topological assumptions—convergence rates. Indeed, regarding applicability, once the rate function is identified, the corresponding inaccuracy rate can be obtained for any new given region of interest. For a detailed exposition of large deviations, we refer to an overview paper [Var08] and the book [DZ09]. As alluded to above, we will use the theory of large deviations to study stochastic policy gradient iterates, that is, to study the iterates of a non-convex optimization algorithm. Although large deviation tools appeared in seminal work on non-convex optimization [GL13], that work does not provide an LDP. What is more, large deviations are intimately connected to rare events, whose analysis is of interest to RL [FMP08]. Also, efficient adaptive sampling techniques become available thanks to careful deployment of large deviation theory [DW04], an approach that is recently attracting interest in the context of SGD [Lah23].
@@
    We find, with high probability, an upper bound on the rate function for iterates generated by softmax policy gradient with an entropy regularized objective (2.3), see Theorem 4.3. We also recover results similar to [MDB20], yet, via large deviation theory, see Lemma 3.2.
@@
    We demonstrate the wide applicability of having established a large deviation rate by leveraging the contraction principle to establish large deviation rates for a wide range of tabular policy parametrizations beyond the softmax parametrization. In particular, we recover the exponential rate from [Mei+20, Thm. 4], yet, as a mere corollary to Theorem 4.3, see Corollary 5.2.
@@
We consider a finite MDP given by a six-tuple $\left(\mathcal{S},\mathcal{A},P,c,\gamma,\rho\right)$ consisting of a finite state space $\mathcal{S}=\{1,\ldots,S\}$, a finite action space $\mathcal{A}=\{1,\ldots,A\}$, a transition kernel $P:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$, a cost-per-stage function $c:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, assumed to be bounded [^1] One can set the cost of a particular stage to be $\infty$ but that results in an infeasible/unbounded optimization problem., a discount factor $\gamma\in(0,1)$, and an initial distribution $\rho\in\Delta(\mathcal{S})$. Here, we use $\Delta(\mathcal{S}):=\{p\in\mathbb{R}^{|\mathcal{S}|}:\ \sum_{i=1}^{|\mathcal{S}|}p_{i}=1,p\geq 0\}$ to denote the probability simplex over $\mathcal{S}$. Throughout the rest of the paper we restrict attention to stationary policies, which are described by a stochastic kernel $\pi\in\Pi=\Delta(\mathcal{A})^{|\mathcal{S}|}$. The value function $V^{\pi}:\Delta(\mathcal{S})\rightarrow\mathbb{R}$ associated with $\pi$ is defined through
$V^{\pi}(\rho):=\mathbb{E}\big{[}\textstyle\sum_{k=0}^{\infty}\gamma^{k}c\left(s_{k},a_{k}\right)|s_{0}\sim\rho,a_{k}\sim\pi\left(\cdot|s_{k}\right),s_{k+1}\sim P(\cdot|s_{k},a_{k})\big{]},$ (2.1)
@@
It is sometimes useful to work with the modified objective where a strongly convex regularizer is added, that is, $V^{\pi}_{\tau}(\rho):=V^{\pi}(\rho)+\tau\cdot H(\rho,\pi)$ [NJG17, Mei+20a], where $\tau>0$ and
$H(\rho,\pi)={\mathbb{E}}\big{[}\textstyle\sum_{k=0}^{\infty}-\gamma^{k}\log\pi\left(a_{k}|s_{k}\right)|s_{0}\sim\rho,a_{k}\sim\pi\left(\cdot|s_{k}\right),s_{k+1}\sim P\left(\cdot|s_{k},a_{k}\right)\big{]}.$
@@
$\pi_{\theta}(a|s)=\frac{\mathrm{exp}\left({\theta(s,a)}\right)}{\sum_{a^{\prime}\in\mathcal{A}}\mathrm{exp}\left({\theta\left(s,a^{\prime}\right)}\right)}\quad\forall a\in\mathcal{A}.$ (2.2)
In the remainder, we frequently think of $\theta$ as a policy parameter in $\mathbb{R}^{d=|\mathcal{S}||\mathcal{A}|}$ and we use $V^{\theta}_{\tau}(\rho)$ for $V^{\pi_{\theta}}_{\tau}(\rho)$. By [Nac+17, Thm. 3] there exists $\theta^{\star}\in\mathbb{R}^{d}$ such that $\pi_{\theta^{\star}}$ is the optimal policy of the entropy-regularized MDP. By the policy gradient theorem [Sut+99], the policy gradient $g(\theta_{t}):=\partial V^{\theta}_{\tau}(\rho)/\partial\theta|_{\theta=\theta_{t}}$ is given by $\partial V^{\theta}_{\tau}(\rho)/\partial\theta(s,a)=1/(1-\gamma)\cdot d_{\rho}^{\theta}(s)\cdot\pi_{\theta}(a|s)\cdot A^{\theta}_{\tau}(s,a)$ , where $d_{\rho}^{\theta}$ denotes the discounted state-visitation frequency measure, and $A^{\theta}_{\tau}$ is the advantage function associated with the regularized objective, see also [Mei+20a, Lem. 1]. In practice, however, computing the advantage function is computationally expensive. It is therefore convenient to work with estimates of the exact gradient. Throughout the rest of the paper we assume unbiased stochastic sample access to $\partial_{\theta}V^{\pi_{\theta}}_{\tau}(\rho)|_{\theta=\theta^{\prime}}$, denoted as $\tilde{g}(\theta^{\prime})$. Stochastic softmax policy gradient with entropy-regularized objective thus suggests to update policy parameters via
$\theta_{t+1}\leftarrow\theta_{t}-\eta_{t}\cdot\tilde{g}(\theta_{t}).$ (2.3)
Assumption 2.1 (Warm start with sufficient exploration).
The initial state distribution satisfies $\min_{s\in\mathcal{S}}\rho(s)>0$. In addition, the initial iterate $\theta_{1}$ is chosen around a $\Delta$-neighborhood of $\theta^{\star}$ for a universal constant $\Delta>0$.
The above assumption simplifies the exposition and such an initial $\theta_{1}$ can be obtained by a stochastic policy gradient method such as [DZL21, Alg. 3.2]. We additionally define $Z_{t}:=g(\theta_{t})-\tilde{g}(\theta_{t}).$ As in [BJK23] we make the following assumption.
Assumption 2.2 (Gradient estimation uncertainty).
@@
By Assumption 2.2 $(i)$, the process $(\theta_{t})_{t\geq 0}$ generated by (2.3) is a Markov chain. We use $\Lambda(\lambda;\theta)=\log\mathbb{E}[\mathrm{exp}(\langle\lambda,Z_{t}\rangle)\mid\theta_{t}=\theta]$ to denote the conditional log-moment generating function (LMGF) of $Z_{t}\in\mathbb{R}^{d}$. We also define $M(\nu;\theta)=\mathbb{E}[\mathrm{exp}(\nu\|Z_{t}\|_{2}^{2})\mid\theta_{t}=\theta]$ as the conditional moment-generating function (MGF) of $\|Z_{t}\|_{2}^{2}$. Now, we impose the following technical assumption to ensure that the tail probabilities of the disturbances $(Z_{t})_{t}$ decay sufficiently fast.
Assumption 2.3 (Sub-Gaussian process).
@@
$\Lambda(\lambda;\theta)\leq\tfrac{1}{2}\sigma^{2}\|\lambda\|_{2}^{2}\quad\forall\lambda,\theta\in\mathbb{R}^{d}.$ (2.4)
A Monte-Carlo gradient estimation method that satisfies the above assumptions is proposed by [DZL21]. The method operates by generating trajectories of MDP with policy parameter $\theta$ and then estimate the advantage function based on these trajectories.
Lemma 2.4 ($L_{1}$-smoothness of the value function [Aga+21, Lem. D.4]).
@@
$|V^{\pi_{\theta}}_{\tau}(\rho)-V^{\pi_{\theta^{\prime}}}_{\tau}(\rho)-\langle g(\theta^{\prime}),\theta-\theta^{\prime}\rangle|\leq\tfrac{1}{2}{L_{1}}\|\theta-\theta^{\prime}\|_{2}^{2}\quad\forall\theta,\theta^{\prime}\in\mathbb{R}^{d}.$ (2.5)
For the dependency of $L_{1}$ on the four-tuple $(c,\gamma,\tau,|\mathcal{A}|)$ see [Mei+20a, Lem. 7 & 14].
@@
If Assumption 2.1 hold, then $\left\|g(\theta)\right\|_{2}^{2}\geq\mu(\theta)\big{(}V_{\tau}^{\theta}(\rho)-V_{\tau}^{\theta^{\star}}(\rho)\big{)},$ where $\mu(\theta)=2\tau|\mathcal{S}|^{-1}\min_{s}\rho(s)$$\min_{s,a}\pi_{\theta}(a|s)^{2}$ $\|d_{\rho}^{\pi_{\theta^{\star}}}/\rho\|_{\infty}^{-1}.$ Moreover, if Assumption 2.2 and 2.3 hold and $\eta_{t}\leq 1/(t+\sqrt{3\sigma^{2}/(2\epsilon\Delta)})$, for iterates generated by (2.3), we have $\mu\!=\inf_{t\geq 1}\mu(\theta_{t})\!>\!0$ with probability at least $61-\epsilon/$ 6.
Proof.
By [Mei+20a, Lem. 15], we have $\left\|g(\theta)\right\|_{2}^{2}\geq\mu(\theta)\big{(}V_{\tau}^{\theta}(\rho)-V_{\tau}^{\theta^{\star}}(\rho)\big{)},$ where $\mu(\theta)=2\tau|\mathcal{S}|^{-1}\min_{s}\rho(s)\min_{s,a}\pi_{\theta}(a|s)^{2}\|d_{\rho}^{\pi_{\theta^{\star}}}/\rho\|_{\infty}^{-1}.$ According to [DZL21, Lem. 6.4] we have further $\inf_{t\geq 1}\pi_{\theta_{t}}(a|s)>0$ with probability $61-\epsilon/$ 6, which concludes the proof. ∎
Lemma 3.1 states that along a trajectory of (2.3), with high probability, we have that $\mu(\theta_{t})>0$ for all $t$. In what follows we will exploit that $\mu\!=\inf_{t\geq 1}\mu(\theta_{t})$ is strictly positive with high probability. Next, we derive elementary recursive inequalities and provide the first main result. If $\eta_{t}$ satisfies $0<\eta_{t}\leq 1/L_{1}$ for all $t=1,\ldots,T$, then (2.3) implies that

### v3

@@ Abstract
Motivated by policy gradient methods in the context of reinforcement learning, we identify a large deviation rate function for the iterates generated by stochastic gradient descent for possibly non-convex objectives satisfying a Polyak-Łojasiewicz condition. Leveraging the contraction principle from large deviations theory, we illustrate the potential of this result by showing how convergence properties of policy gradient with a softmax parametrization and an entropy regularized objective can be naturally extended to a wide spectrum of other policy parametrizations.
@@
Policy gradient methods are at the core of several reinforcement learning (RL) algorithms. e.g., see [Sutton et al. (1999), Degris et al. (2012), Agarwal et al. (2021)]. As such, a wide adoption of these algorithms is aided by precisely understanding their performance. To that end, in scenarios where policy gradients can only be evaluated stochastically, it is crucial to understand the underlying distribution governing the evolution of policy iterates. Despite the popularity of stochastic policy gradient methods for over 20 years, global convergence has been only understood recently (Fazel et al., 2018; Zhang et al., 2020; Bhandari and Russo, 2019; Agarwal et al., 2021). Although global convergence proofs are actively developed for different policy gradient algorithm templates (Bhandari and Russo, 2019; Agarwal et al., 2021; Cen et al., 2022), two aspects of the current convergence analysis would benefit from further study: 1) global convergence guarantees are often stated in terms of expected suboptimality; and 2) the choice of policy parametrizations impacts convergence behavior (Mei et al., 2020a). In light of these observations, this paper aims to derive sharp convergence rates in probability and to provide a unifying approach towards understanding the effect of different policy parametrizations. As we build upon a rich history of work, we succinctly comment on: (i) high probability analysis in stochastic optimization and RL; and (ii) policy gradient methods in RL.
@@
(i) Despite stochastic gradient descent (SGD) being over 70 years old (Robbins and Monro, 1951), there has been a surge of interest in terms of high probability analysis for (SGD): for strongly-convex objectives (Harvey et al., 2019); non-convex objectives satisfying a Polyak-Łojasiewicz (PL) condition (Madden et al., 2020); and non-convex objectives (Ghadimi and Lan, 2013; Liu et al., 2023) with Lipschitz gradients. A motivation for studying tight convergence bounds in terms of probabilities over convergence in expectation is due to the fact that the iterates generated by SGD can be brittle and large deviations from the expected value can occur, e.g., see Gower et al. (2020) and references therein. Hence, a high-probability bound is attractive since it reassures the practitioner that certain behavior occurs at least with a close-to-$1$ probability for a single realization of iterates.
@@
(ii) Policy gradient methods can be interpreted as gradient descent applied to the policy optimization problem in RL (Sutton and Barto, 2018, Ch. 13), see in particular (Sutton and Barto, 2018, p. 337) for historical remarks. Therefore, when the state space is large (Bottou, 2010) and due to necessary approximation, one oftentimes resorts to stochastic variants, whose convergence guarantees are often stated in terms of expectations (Lan, 2023), with the exception of (Ding et al., 2021), see also (Madden et al., 2020, § 2.2). The high-probability convergence rate provided in Ding et al. (2021) can be seen as a loosened bound of the large deviation rate we aim to derive. The pursuit of a tight high-probability concentration bound for the iterations produced by stochastic policy gradient methods holds significant potential for bolstering the practical implementation and interpretability of policy gradient methods, thereby improving the overall applicability of RL. Studying the concentration behavior of SGD iterates through the lens of large deviations is pioneered by Bajovic et al. (2023). It is worth mentioning that, however, their analysis exploits strong convexity of the optimization objective. Then, a first step towards understanding more precisely the concentration behavior of stochastic policy gradient iterates would be to study its entropy-regularized variant, which already faces the difficulties of non-convexity and non-uniformity of the PL constant.
@@
As the theory of large deviations is the key tool in this paper, we briefly introduce it already at this stage. Let $X_{t}\in\mathbb{R}^{d}$ be the $t^{\mathrm{th}}$ iterate of some stochastic algorithm aimed at driving $X_{t}$ to $x^{\star}$ for $t\to+\infty$, then, many aforementioned works provide statistical guarantees of the form $\mathbb{P}(\|X_{t}-x^{\star}\|_{2}\leq\varepsilon)\geq 1-\beta$ $\forall t\geq T$ for an appropriately chosen triple $(T,\varepsilon,\beta)$. Although often impressive pieces of work, there are some remarks to be made. First, it is often not clear if these high-probability bounds are tight. Secondly, the discrepancy is often measured by a sufficiently simple function, like the $\ell_{2}$-norm in this case, however, in practice one might be interested in vastly different sublevel sets, e.g., sublevel sets capturing harmful events, the ones that are of interest in safe RL. Third, the guarantees are oftentimes states for particular policy parametrizations, a unifying framework that generalizes performance analysis across different policy parametrizations remains largely unexplored. The theory of large deviations is precisely powerful when one aims to address these aforementioned points. Let us clarify terminology and provide intuition, yet, while skipping some details. A sequence of finite (probability) measures $(\mu_{t})_{t}$ on $\mathbb{R}^{d}$ is said to satisfy a large deviation principle (LDP) with rate function $I:\mathbb{R}^{d}\to[0,+\infty]$ when for any Borel set $\Theta\subseteq\mathbb{R}^{d}$
$-\mathring{r}:=-\inf_{\theta\in\mathring{\Theta}}I(\theta)\leq\liminf_{t\to+\infty}\frac{1}{t}\log\mu_{t}({\Theta})\leq\limsup_{t\to+\infty}\frac{1}{t}\log\mu_{t}({\Theta})\leq-\inf_{\theta\in\overline{\Theta}}I(\theta)=:-\overline{r},$ (1)
@@
where $\mathring{\Theta}$ denotes the interior of $\Theta$, and $\overline{\Theta}$ denotes the closure of $\Theta.$ Now, identifying $(\mu_{t})_{t}$ with a sequence of random variables $(X_{t})_{t}$, we have after rearranging (1) that $\mathrm{exp}(-\mathring{r}t+o(t))\leq\mathbb{P}(X_{t}\in\Theta)\leq\mathrm{exp}(-\overline{r}t+o(t)).$ As such, an LDP captures—possibly tight, e.g., under mild topological assumptions—convergence rates. Indeed, regarding applicability, once the rate function is identified, the corresponding inaccuracy rate can be obtained for any new given region of interest. For a detailed exposition of large deviations, we refer to an overview paper by Varadhan (2008) and the book by Dembo and Zeitouni (2009). As alluded to above, we will use the theory of large deviations to study stochastic policy gradient iterates, that is, to study the iterates of a non-convex optimization algorithm. Although large deviation tools appeared in seminal work on non-convex optimization Ghadimi and Lan (2013), that work does not provide an LDP. What is more, large deviations are intimately connected to rare events, whose analysis is of interest to RL (Frank et al., 2008). Also, efficient adaptive sampling techniques become available thanks to careful deployment of large deviation theory (Dupuis and Wang, 2004), an approach that is recently attracting interest in the context of SGD (Lahire, 2023).
@@
    We find, with high probability, an lower bound on the rate function for iterates generated by softmax policy gradient with an entropy regularized objective (4), see Theorem 4.3. We also recover results similar to Madden et al. (2020), yet, via large deviation theory, see Lemma 3.3.
@@
    We demonstrate the wide applicability of having established a large deviation rate by leveraging the contraction principle to establish large deviation rates for a wide range of tabular policy parametrizations beyond the softmax parametrization.
@@
We consider a finite MDP given by a six-tuple $\left(\mathcal{S},\mathcal{A},P,c,\gamma,\rho\right)$ consisting of a finite state space $\mathcal{S}=\{1,\ldots,S\}$, a finite action space $\mathcal{A}=\{1,\ldots,A\}$, a transition kernel $P:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{P}(\mathcal{S})$, a cost-per-stage function $c:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, assumed to be bounded [^1] One can set the cost of a particular stage to be $\infty$ but that results in an infeasible/unbounded optimization problem., a discount factor $\gamma\in(0,1)$, and an initial distribution $\rho\in\mathcal{P}(\mathcal{S})$. Here, we use $\mathcal{P}(\mathcal{S}):=\{p\in\mathbb{R}^{|\mathcal{S}|}:\ \sum_{i=1}^{|\mathcal{S}|}p_{i}=1,p\geq 0\}$ to denote the probability simplex over $\mathcal{S}$. Throughout the rest of the paper we restrict attention to stationary policies, which are described by a stochastic kernel $\pi\in\Pi:=\mathcal{P}(\mathcal{A})^{|\mathcal{S}|}$. The value function $V^{\pi}:\mathcal{P}(\mathcal{S})\rightarrow\mathbb{R}$ associated with $\pi$ is defined through
$V^{\pi}(\rho):=\mathbb{E}\big{[}\textstyle\sum_{k=0}^{\infty}\gamma^{k}c\left(s_{k},a_{k}\right)|s_{0}\sim\rho,a_{k}\sim\pi\left(\cdot|s_{k}\right),s_{k+1}\sim P(\cdot|s_{k},a_{k})\big{]},$ (2)
@@
It is sometimes useful to work with the modified objective where a regularizer is added, that is, $V^{\pi}_{\tau}(\rho):=V^{\pi}(\rho)+\tau\cdot\mathbb{H}(\rho,\pi)$ (Neu et al., 2017; Mei et al., 2020b), where $\tau>0$ and
$\mathbb{H}(\rho,\pi):={\mathbb{E}}\big{[}\textstyle\sum_{k=0}^{\infty}-\gamma^{k}\log\pi\left(a_{k}|s_{k}\right)|s_{0}\sim\rho,a_{k}\sim\pi\left(\cdot|s_{k}\right),s_{k+1}\sim P\left(\cdot|s_{k},a_{k}\right)\big{]}.$
@@
$\pi_{\theta}(a|s)=\frac{\mathrm{exp}\left({\theta(s,a)}\right)}{\sum_{a^{\prime}\in\mathcal{A}}\mathrm{exp}\left({\theta\left(s,a^{\prime}\right)}\right)}\quad\forall a\in\mathcal{A}.$ (3)
In the remainder, we frequently think of $\theta$ as a policy parameter in $\mathbb{R}^{d=|\mathcal{S}||\mathcal{A}|}$ and we use $V^{\theta}_{\tau}(\rho)$ for $V^{\pi_{\theta}}_{\tau}(\rho)$. By (Nachum et al., 2017, Thm. 3) there exists $\theta^{\star}\in\mathbb{R}^{d}$ such that $\pi_{\theta^{\star}}$ is the optimal policy of the entropy-regularized MDP. By the policy gradient theorem (Sutton et al., 1999), the policy gradient $g(\theta_{t}):=\partial V^{\theta}_{\tau}(\rho)/\partial\theta|_{\theta=\theta_{t}}$ is given by $\partial V^{\theta}_{\tau}(\rho)/\partial\theta(s,a)=1/(1-\gamma)\cdot d_{\rho}^{\theta}(s)\cdot\pi_{\theta}(a|s)\cdot A^{\theta}_{\tau}(s,a)$ , where $d_{\rho}^{\theta}$ denotes the discounted state-visitation frequency measure, and $A^{\theta}_{\tau}$ is the advantage function associated with the regularized objective, see also (Mei et al., 2020b, Lem. 1). In practice, however, computing the advantage function is computationally expensive. It is therefore convenient to work with estimates of the exact gradient. Throughout the rest of the paper we assume unbiased stochastic sample access to $\partial_{\theta}V^{\pi_{\theta}}_{\tau}(\rho)|_{\theta=\theta^{\prime}}$, denoted as $\tilde{g}(\theta^{\prime})$. Stochastic softmax policy gradient with entropy-regularized objective thus suggests to update policy parameters via
$\theta_{t+1}\leftarrow\theta_{t}-\eta_{t}\cdot\tilde{g}(\theta_{t}).$ (4)
Assumption 1 (Warm start with sufficient exploration)
The initial state distribution satisfies $\allowbreak\min_{s\in\mathcal{S}}\rho(s)>0$. In addition, the initial iterate $\theta_{1}$ is chosen around a $\Delta$-neighborhood of $\theta^{\star}$ in the sense that $\min_{\theta^{\star}\in\Theta^{\star}}\|\theta_{1}-\theta^{\star}\|_{2}\leq\Delta$ where $\Theta^{\star}$ is the set of all optimal solutions.
The above assumption simplifies the exposition and such an initial $\theta_{1}$ can be obtained by a stochastic policy gradient method such as (Ding et al., 2021, Alg. 3.2). We additionally define $Z_{t}:=g(\theta_{t})-\tilde{g}(\theta_{t}).$ As in (Bajovic et al., 2023) we make the following assumption.
Assumption 2 (Gradient estimation uncertainty)
@@
By Assumption 2 $(i)$, the process $(\theta_{t})_{t\geq 0}$ generated by (4) is a Markov chain. We use $\Lambda(\lambda;\theta)=\log\mathbb{E}[\mathrm{exp}(\langle\lambda,Z_{t}\rangle)\mid\theta_{t}=\theta]$ to denote the conditional log-moment generating function (LMGF) of $Z_{t}\in\mathbb{R}^{d}$. We also define $M(\nu;\theta)=\mathbb{E}[\mathrm{exp}(\nu\|Z_{t}\|_{2}^{2})\mid\theta_{t}=\theta]$ as the conditional moment-generating function (MGF) of $\|Z_{t}\|_{2}^{2}$. Now, we impose the following technical assumption to ensure that the tail probabilities of the disturbances $(Z_{t})_{t}$ decay sufficiently fast.
Assumption 3 (Sub-Gaussian process)
@@
$\Lambda(\lambda;\theta)\leq\tfrac{1}{2}\sigma^{2}\|\lambda\|_{2}^{2}\quad\forall\lambda,\theta\in\mathbb{R}^{d}.$ (5)
A Monte-Carlo gradient estimation method that satisfies the above assumptions is proposed by Ding et al. (2021). The method operates by generating trajectories of the MDP with policy parameter $\theta$ and then estimate the advantage function based on these trajectories.
Lemma 2.1 ($L_{1}$-smoothness of the value function (Mei et al., 2020b, Lem. 7 & 14)).
@@
$|V^{\pi_{\theta}}_{\tau}(\rho)-V^{\pi_{\theta^{\prime}}}_{\tau}(\rho)-\langle g(\theta^{\prime}),\theta-\theta^{\prime}\rangle|\leq\tfrac{1}{2}{L_{1}}\|\theta-\theta^{\prime}\|_{2}^{2}\quad\forall\theta,\theta^{\prime}\in\mathbb{R}^{d}.$ (6)
For the dependency of $L_{1}$ on the four-tuple $(c,\gamma,\tau,|\mathcal{A}|)$ see (Mei et al., 2020b, Lem. 7 & 14).
@@
If Assumption 1 holds, then $\left\|g(\theta)\right\|_{2}^{2}\geq\mu(\theta)\big{(}V_{\tau}^{\theta}(\rho)-V_{\tau}^{\theta^{\star}}(\rho)\big{)},$ where $\mu(\theta)=2\tau|\mathcal{S}|^{-1}\min_{s}\rho(s)\min_{s,a}\pi_{\theta}(a|s)^{2}\|d_{\rho}^{\pi_{\theta^{\star}}}/\rho\|_{\infty}^{-1}.$ Moreover, if Assumption 2 and 3 hold and $\eta_{t}\leq\eta/(t+\sqrt{3\sigma^{2}/(2\epsilon\Delta)})$, for iterates generated by (4), we have $\mu\!=\inf_{t\geq 1}\mu(\theta_{t})\!>\!0$ with probability at least $61-\epsilon/$ 6.
Proof 3.2.
By (Mei et al., 2020b, Lem. 15), we have $\left\|g(\theta)\right\|_{2}^{2}\geq\mu(\theta)\big{(}V_{\tau}^{\theta}(\rho)-V_{\tau}^{\theta^{\star}}(\rho)\big{)},$ where $\mu(\theta)=2\tau|\mathcal{S}|^{-1}\min_{s}\rho(s)\min_{s,a}\pi_{\theta}(a|s)^{2}\|d_{\rho}^{\pi_{\theta^{\star}}}/\rho\|_{\infty}^{-1}.$ According to (Ding et al., 2021, Lem. 6.4) we further have $\inf_{t\geq 1}\pi_{\theta_{t}}(a|s)>0$ with probability $61-\epsilon/$ 6, which concludes the proof.
Lemma 3.1 states that along a trajectory of (4), with high probability, we have that $\mu(\theta_{t})>0$ for all $t$. In what follows we will exploit that $\mu\!=\inf_{t\geq 1}\mu(\theta_{t})$ is strictly positive with high probability. Next, we derive elementary recursive inequalities and provide the first main result. If $\eta_{t}$ satisfies $0<\eta_{t}\leq 1/L_{1}$ for all $t=1,\ldots,T$, then (4) implies that